# ROVER: A Multi-Season Dataset for Visual SLAM

### v1

@@ V-D Discussion
Seasonal changes introduce substantial variability in SLAM performance, particularly in summer and autumn when denser vegetation and shifting lighting make feature extraction difficult. Algorithms such as ORB-SLAM3 and DROID-SLAM perform better during winter and spring but experience notable declines in accuracy during summer (refer to Table X). These seasonal patterns suggest that current SLAM systems are not yet optimized for environments where visual complexity fluctuates due to natural factors such as vegetation growth.
Lighting conditions further exacerbate these challenges. While SLAM systems perform well in daylight, as evidenced by the low ATEs and high success rates, performance sharply declines in low-light scenarios such as dusk and night. Even with external lighting, many methods still struggle to maintain accuracy, as shown by the increase in ATE values at night (refer to Table XI). This indicates that the feature extraction process in low-light environments remains a significant bottleneck for SLAM systems. Qualitatively, this is evident in locations like Campus Large and Park, where even with controlled environments, systems like SVO Pro exhibit drift and jitter under suboptimal lighting, contributing to higher error rates.
The camera and sensor configurations also play a crucial role in determining SLAM accuracy. Mono cameras, for instance, tend to perform adequately in smaller environments such as Garden Small but falter in larger, more complex settings. Monocular approaches, such as DPVO and DPV-SLAM, heavily struggle with scaling issues due to the inherent scale ambiguity of these systems, particularly in larger locations such as Campus Large and Garden Large, where visual cues are more difficult to anchor (refer to Figure 7). In contrast, RGB-D configurations such as ORB-SLAM3 (RGB-D D435i) benefit from depth data, leading to improved performance in both small and large environments. However, adding inertial data does not consistently improve performance. In some cases, as seen with RGB-D-inertial setups, inertial data even reduces accuracy and success rates, suggesting that the integration of visual and inertial data requires further refinement.
The results of this evaluation underscore the need for more robust SLAM systems capable of adapting to the unpredictable dynamics of outdoor environments. While some configurations, such as OpenVINS (stereo-inertial T265 internal) and DROID-SLAM (RGB-D D435i), show promising results, many algorithms still struggle with low-light conditions, high vegetation density, and large, unstructured environments. This dataset reveals critical challenges that SLAM systems must overcome to achieve long-term, robust localization and mapping in natural outdoor settings. Further research is needed to enhance the resilience of the SLAM algorithms to environmental variability and improve their adaptability across diverse scenarios.
@@ VI Conclusion
In this paper, we presented a large-scale benchmark dataset designed to evaluate SLAM algorithms under a variety of spatial and environmental conditions, allowing the assessment of complex real-world scenarios. The benchmark results demonstrate that while some SLAM configurations, particularly those using RGB-D or stereo-inertial setups, show promising performance in structured environments, challenges remain significant under varying seasonal and illumination conditions. Notably, performance degrades in low-light and high-vegetation settings, with substantial increases in trajectory errors during summer and autumn, where denser vegetation and variable lighting complicate feature extraction.
Lighting conditions further impact robustness, as most SLAM methods achieve reliable results in daylight but experience sharp declines in accuracy at dusk and night. Despite improvements with external lighting, many configurations, especially monocular approaches, continue to struggle with maintaining accuracy due to scale ambiguity and drift. The qualitative analysis underscores that algorithms such as ORB-SLAM3 and DROID-SLAM show resilience in structured or well-lit conditions but face limitations in environments with limited visual features or uneven lighting, such as rooftop spaces or shaded garden areas. These results highlight that while current SLAM methods have made strides, their robustness in highly variable natural environments is limited, especially regarding long-term mapping and consistency. Further development is needed to enhance SLAM resilience in outdoor settings, and this dataset provides a comprehensive foundation for addressing these gaps.

### v2

@@ V-D Discussion
Seasonal changes introduce substantial variability in SLAM performance, particularly in summer and autumn when denser vegetation and shifting lighting make feature extraction difficult. Algorithms such as ORB-SLAM3 and DROID-SLAM perform better during winter and spring but experience notable declines in accuracy during summer (refer to Table XI). These seasonal patterns suggest that current SLAM systems are not yet optimized for environments where visual complexity fluctuates due to natural factors such as vegetation growth.
Lighting conditions further exacerbate these challenges. While SLAM systems perform well in daylight, as evidenced by the low mATEs, low mRPEs and high SR, performance sharply declines in low-light scenarios such as dusk and night. Even with external lighting, many methods still struggle to maintain accuracy, as shown by the increase in mATE and mRPE values at night (refer to Table XII). This indicates that the feature extraction process in low-light environments remains a significant bottleneck for SLAM systems. Qualitatively, this is evident in locations like Campus Large and Park, where even with controlled environments, systems like SVO Pro exhibit drift and jitter under suboptimal lighting, contributing to higher error rates.
The camera and sensor configurations also play a crucial role in determining SLAM accuracy. Mono cameras, for instance, tend to perform adequately in smaller environments such as Garden Small but falter in larger, more complex settings. Monocular approaches, such as DPVO and DPV-SLAM, heavily struggle with scaling issues due to the inherent scale ambiguity of these systems, particularly in larger locations such as Campus Large and Garden Large, where visual cues are more difficult to anchor (refer to Figure 9). In contrast, RGBD configurations such as ORB-SLAM3 (RGBD D435i) benefit from depth data, leading to improved performance in both small and large environments. However, adding inertial data does not consistently improve performance. In some cases, as seen with RGBD-inertial setups, inertial data even reduces accuracy and success rates, suggesting that the integration of visual and inertial data requires further refinement.
The results of this evaluation underscore the need for more robust SLAM systems capable of adapting to the unpredictable dynamics of outdoor environments. While some configurations, such as OpenVINS (stereo-inertial T265 internal) and DROID-SLAM (RGBD D435i), show promising results, many algorithms still struggle with low-light conditions, high vegetation density, and large, unstructured environments. This dataset reveals critical challenges that SLAM systems must overcome to achieve long-term, robust localization and mapping in natural outdoor settings. Further research is needed to enhance the resilience of the SLAM algorithms to environmental variability and improve their adaptability across diverse scenarios.
@@ VI Conclusion
In this paper, we presented a large-scale benchmark dataset designed to evaluate SLAM algorithms under a variety of spatial and environmental conditions, allowing the assessment of complex real-world scenarios. The benchmark results demonstrate that while some SLAM configurations, particularly those using RGBD or stereo-inertial setups, show promising performance in structured environments, challenges remain significant under varying seasonal and lighting conditions. Notably, performance degrades in low-light and high-vegetation settings, with substantial increases in trajectory errors during summer and autumn, where denser vegetation and variable lighting complicate feature extraction.
Lighting conditions further impact robustness, as most SLAM methods achieve reliable results in daylight but experience sharp declines in accuracy at dusk and night. Despite improvements with external lighting, many configurations, especially monocular approaches, continue to struggle with maintaining accuracy due to scale ambiguity and drift. The qualitative analysis underscores that algorithms such as ORB-SLAM3 and DROID-SLAM show resilience in structured or well-lit conditions but face limitations in environments with limited visual features or uneven lighting, such as rooftop spaces or shaded garden areas. These results highlight that while current SLAM methods have made strides, their robustness in highly variable semi-structured natural environments is limited, especially regarding long-term mapping and consistency. Further development is needed to enhance SLAM resilience in outdoor settings, and this dataset provides a comprehensive foundation for addressing these gaps.